# Language Supervised Human Action Recognition with Salient Fusion: Construction Worker Action Recognition as a Use-Case

### v1

@@ V-D Ablation Studies
In this section, we examine the performance of each module using the NTU-RGB+D dataset, summarized in Table III. We first evaluate the efficacy of individual modalities (skeleton or visual) for HAR, both with and without a text encoder. Our analysis shows that the visual modality performs better, while the text encoder enhances the performance of the skeleton encoder. Notably, merging the two modalities with our fusion module shows that using text supervision for both or neither modality decreases performance by 0.5% and 0.6% on XSub benchmarks compared to using it only for the skeleton encoder. This is due to the superior performance of our visual module and the fusion module’s effectiveness, making text supervision unnecessary for the visual encoder.
In another aspect of our study, we assessed our fusion module without the salient feature extraction, relying instead on self-attention for each modality’s features. Our ablation study revealed that extracting salient features improves model performance by 0.8%. Additionally, learnable prompts enhance performance by 0.3%, and conditioning on skeleton features adds another 0.1%. For comparison, using the ViT-B/32 image encoder resulted in a 2.3% performance decrease.We conducted similar ablations on the VolvoConstAct dataset as performed on the NTU-RGB+D dataset, and the results were consistent. We established a baseline for action recognition of construction workers using our dataset. Enhanced performance for the skeleton modality on VolvoConstAct could be achieved with additional denoising and pre-processing, as we used raw skeleton data to test the model’s robustness and establish our evaluations as a baseline. Additionally, our high performance suggests that future studies could benefit from focusing solely on the visual modality, which is more advantageous for long-range HAR.
TABLE III: Ablation study for our method on the NTU-RGB+D [54] (NTU60) and VolvoConstAct (Volvo) datasets.
Method NTU60 (%) Volvo
XSub XView (%)
Only Skeletons 86.2 92.5 15.2
Only Skeletons Supervised with Text 86.6 94.5 15.5
Only Video 91.2 96.5 92.9
Only Video Supervised with Text 91.3 96.0 92.9
Fusion without fine-grained down-sampling 92.9 97.0 94.3
Ours without Text Supervision on Skeletons or Video 93.1 96.7 95.2
Ours with Supervising both Skeletons and Video 93.2 96.5 96.4
Ours with Non-learnable Prompts (Pre-defined Text Prompts) 93.4 97.1 93.6
Ours without Conditional Prompts (Without Adding Extracted Skeleton Features to the Learnable Prompts) 93.6 97.4 94.5
Ours with ViT-B/32 91.4 95.7 91.5
Ours 93.7 97.3 96.4

### v2

@@ V-D Ablation Studies
In this section, we examine the performance of each module using the NTU-RGB+D dataset, summarized in Table IV. We first evaluate the efficacy of individual modalities (skeleton or visual) for HAR, both with and without a text encoder. Our analysis shows that the visual modality performs better, while the text encoder enhances the performance of the skeleton encoder. Notably, merging the two modalities with our fusion module shows that using text supervision for both or neither modality decreases performance by 0.5% and 0.6% on XSub benchmarks compared to using it only for the skeleton encoder. This is due to the superior performance of our visual module and the fusion module’s effectiveness, making text supervision unnecessary for the visual encoder.
In another aspect of our study, we assessed our fusion module without the salient feature extraction, relying instead on self-attention for each modality’s features. Our ablation study revealed that extracting salient features improves model performance by 0.8%. Additionally, learnable prompts enhance performance by 0.3%, and conditioning on skeleton features adds another 0.1%, both of which remain valuable given our competitive benchmark. For comparison, using the ViT-B/32 image encoder resulted in a 2.3% performance decrease.
TABLE IV: Ablation study for our method on the NTU-RGB+D [55] (NTU60) dataset.
Method NTU60 (%)
XSub XView
Only Skeletons 86.2 92.5
Only Skeletons Supervised with Text 86.6 94.5
Only Video 91.2 96.5
Only Video Supervised with Text 91.3 96.0
Fusion without fine-grained down-sampling 92.9 97.0
Ours without Text Supervision on Skeletons or Video 93.1 96.7
Ours with Supervising both Skeletons and Video 93.2 96.5
Ours with Non-learnable Prompts (Pre-defined Text Prompts) 93.4 97.1
Ours without Conditional Prompts (Without Adding Extracted Skeleton Features to the Learnable Prompts) 93.6 97.4
Ours with ViT-B/32 91.4 95.7
Ours 93.7 97.3
@@ VI Conclusion
ACKNOWLEDGMENT
We sincerely acknowledge Mr. Mikael Fries for his contributions to real machine development and his insightful discussions.